# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ H.1 Proof of Technical Lemmas
$\delta_{0}=\kappa_{0}^{\beta}M^{-\beta}.$ (H.28)
Under the assumption that $d_{\texttt{non}}$ is a fixed constant in Lemma H.5, we have $\kappa_{0}$ is also a constant and $\delta_{0}=\Theta(M^{-\beta})$. Under Equation (H.28), the bound in Equation (H.26) becomes
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[(f_{\bm{\omega}}(\bm{O}_{t})-f_{\bm{\omega}^{\prime}}(\bm{O}_{t}))^{2}\mathbbm{1}\left\{A_{t}=1,V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right]$ (H.29)
$~~\lesssim M^{-2\beta-d_{\texttt{non}}}+M^{-2\beta}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
and Equation (H.27) becomes
$\mathbb{E}_{f_{\bm{\omega}}}\left[\left(f_{\bm{\omega}}\left(\bm{O}_{i}^{(0)}\right)-f_{\bm{\omega}^{\prime}}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}\right]\lesssim M^{-2\beta-d_{\texttt{non}}}.$ (H.30)
Combining Equations (H.29) and (H.30) with Equation (H.1), for any $t\in[T]$ and the choice of $\delta_{0}$ given in Equation (H.28),
$\mathcal{K}_{\bm{\theta},t}^{(\texttt{non})}\left(f_{\bm{\omega}},f_{\bm{\omega}^{\prime}}\right)$ $\lesssim M^{-2\beta}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}=1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]+\frac{(t+N)}{M^{2\beta+d_{\texttt{non}}}}.$ (H.31)
Using an average hammer over $\bm{\omega}\in\Omega_{m}$, it follows from Equation (H.24) and the choice of $\delta_{0}$ in Equation (H.28) that
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)$ $\gtrsim M^{-\beta}\sup_{\bm{\omega}\in\Omega_{m}}\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$ (H.32)
$\geq 2^{-m}M^{-\beta}\sum_{\bm{\omega}\in\Omega_{m}}\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right],$
@@
$G_{j}^{t}:=\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{i\in\{\pm 1\}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq i,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right],$ (H.33)
where we group $\bm{\omega}_{[-j]}^{1}$ and $\bm{\omega}_{[-j]}^{-1}$ together in the inner sum. Taking Equation (H.33) into Equation (H.32), we have
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)\gtrsim 2^{-m}M^{-\beta}\sum_{j=1}^{m}\sum_{t=1}^{T}G_{j}^{t}.$ (H.34)
We pause to provide some intuition for introducing $G_{j}^{t}$. The idea is that we would like to apply Bretagnolle-Huber inequality as stated in Theorem H.1 to obtain a lower bound of the cumulative regret. To get a tighter lower bound, we would group the most similar pairs of $\bm{\omega},\bm{\omega}^{\prime}\in\Omega_{m}$ together to minimize the KL divergence between the two probability measures indexed by $\bm{\omega}$ and $\bm{\omega}^{\prime}$.
By Equation (H.25) and the definition of $\delta_{0}$ in Equation (H.28),
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq i,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$ $\asymp\frac{1}{M^{d_{\texttt{non}}}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq i\right\}\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right].$ (H.35)
Denote by $\mathbb{P}^{(t-1)}_{j}$ the conditional probability $\mathbb{P}\left(\cdot\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right)$. We apply Bretagnolle-Huber inequality as stated in Theorem H.1 and obtain
$\sum_{i\in\{\pm 1\}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{t}\neq i\right\}\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right]$ (H.36)
@@
where the last inequality follows from Equations (H.16), (H.1) and (H.2). Taking Equations (H.35) and (H.36) into Equation (H.33) yields that
$G_{j}^{t}$ $\gtrsim M^{-d_{\texttt{non}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left[-\mathcal{K}_{\bm{\theta},t}^{(\texttt{non})}\left(f_{\bm{\omega}^{1}_{[-j]}},f_{\bm{\omega}^{-1}_{[-j]}}\right)\right]$ (H.37)
$\stackrel{{\scriptstyle(i)}}{{\geq}}\frac{1}{M^{d_{\texttt{non}}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left(-\frac{C}{M^{2\beta}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(t+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
$\geq\frac{1}{M^{d_{\texttt{non}}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left(-\frac{C}{M^{2\beta}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(T+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}\frac{2^{m-1}}{M^{d_{\texttt{non}}}}\exp\left(-\frac{C}{M^{2\beta}2^{m-1}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(T+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
where inequality $(i)$ follows from Equation (H.31) and inequality $(ii)$ follows from Jensen’s inequality. Let
$E_{j,\pi}:=\frac{1}{2^{m-1}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathbbm{1}\left\{A_{\tau}=1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]$
and taking $E_{j,\pi}$ into Equation (H.37), we have
$G_{j}^{t}\gtrsim\frac{2^{m-1}}{M^{d_{\texttt{non}}}}\exp\left(-CM^{-2\beta}E_{j,\pi}-C(T+N)M^{-2\beta-d_{\texttt{non}}}\right)$ (H.38)
From the definition of $G_{j}^{t}$ in Equation (H.33), we also have
$\sum_{t=1}^{T}G_{j}^{t}\geq 2^{m-1}E_{j,\pi}.$ (H.39)
Taking Equations (H.38) and (H.39) into Equation (H.34) yields
@@
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)$ $\gtrsim\inf_{z\geq 0}mM^{-\beta}\left\{z+\frac{T}{M^{d_{\texttt{non}}}}\exp\left[-CM^{-2\beta}z-\frac{CN}{M^{2\beta+d_{\texttt{non}}}}\right]\right\}$ (H.40)
@@
where in the last inequality, we use the definition of $m$ as in Equation (H.9) and let
@@
The minimizer of the right-hand side of Equation (H.40) over $z\in{{\mathbb{R}}}$ is given by
$z^{\star}=\frac{M^{2\beta}}{C}\log\left(\frac{CT\alpha}{M^{2\beta+d_{\texttt{non}}}}\right)=\frac{M^{2\beta}}{C}\log\left(\frac{CT}{M^{2\beta+d_{\texttt{non}}}}\right)-\frac{N}{M^{d_{\texttt{non}}}}.$ (H.41)
@@
$M^{2\beta+d_{\texttt{non}}}\log\left(\frac{CT}{M^{2\beta+d_{\texttt{non}}}}\right)\geq CN.$ (H.42)
Noting that when $M^{2\beta+d_{\texttt{non}}}>CT$, the left hand side of the above display is negative. Thus, for Equation (H.42) to hold, we must have $M^{2\beta+d_{\texttt{non}}}=O(T)$, implying that
@@
$M^{2\beta+d_{\texttt{non}}}=\frac{CT}{e}.$ (H.43)
When $T\geq CN$ for some constant $C$ sufficiently large, $z^{\star}\geq 0$ holds. Taking $z=z^{\star}$ in Equation (H.40) yields
@@
where the last equality holds from the choice of $M$ in Equation (H.43).
@@
and is monotonically increasing when $z>z_{0}$, it follows that the minimizer of $h(z)$ over $z\geq 0$ when $z_{0}<0$ is attained at $z=0$. Comparing the form of the right-hand side of Equation (H.40) to $h(z)$ defined above yields that the minimizer is attained at $z=0$ and
$\sup_{f}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)\geq\frac{T}{M^{\beta}}\exp\left(-\frac{CN}{M^{2\beta+d_{\texttt{non}}}}\right).$ (H.44)
@@
$M=\left[\frac{C(2\beta+d_{\texttt{non}})N}{\beta}\right]^{\frac{1}{2\beta+d_{\texttt{non}}}}=\Theta\left(N^{\frac{1}{2\beta+d_{\texttt{non}}}}\right).$ (H.45)
Taking Equation (H.45) into the right-hand side of Equation (H.44) yields the desired result in Equation (H.13). ∎
Appendix I Derivation of Equivalent Formulations for UCB Exploration in Algorithm 1

### v3

@@ I.2 Theoretical Guarantees and Regret Comparison
$\delta_{0}=\kappa_{0}^{\beta}M^{-\beta}.$ (K.28)
Under the assumption that $d_{\texttt{non}}$ is a fixed constant in Lemma K.5, we have $\kappa_{0}$ is also a constant and $\delta_{0}=\Theta(M^{-\beta})$. Under Equation (K.28), the bound in Equation (K.26) becomes
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[(f_{\bm{\omega}}(\bm{O}_{t})-f_{\bm{\omega}^{\prime}}(\bm{O}_{t}))^{2}\mathds{1}\left\{A_{t}=1,V_{t}=0\right\}\mid\mathcal{H}_{t-1}\right]$ (K.29)
$~~\lesssim M^{-2\beta-d_{\texttt{non}}}+M^{-2\beta}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}=1,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$
and Equation (K.27) becomes
$\mathbb{E}_{f_{\bm{\omega}}}\left[\left(f_{\bm{\omega}}\left(\bm{O}_{i}^{(0)}\right)-f_{\bm{\omega}^{\prime}}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}\right]\lesssim M^{-2\beta-d_{\texttt{non}}}.$ (K.30)
Combining Equations (K.29) and (K.30) with Equation (K.1), for any $t\in[T]$ and the choice of $\delta_{0}$ given in Equation (K.28),
$\mathcal{K}_{\bm{\theta},t}^{(\texttt{non})}\left(f_{\bm{\omega}},f_{\bm{\omega}^{\prime}}\right)$ $\lesssim M^{-2\beta}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}=1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]+\frac{(t+N)}{M^{2\beta+d_{\texttt{non}}}}.$ (K.31)
Using an average hammer over $\bm{\omega}\in\Omega_{m}$, it follows from Equation (K.24) and the choice of $\delta_{0}$ in Equation (K.28) that
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)$ $\gtrsim M^{-\beta}\sup_{\bm{\omega}\in\Omega_{m}}\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$ (K.32)
$\geq 2^{-m}M^{-\beta}\sum_{\bm{\omega}\in\Omega_{m}}\sum_{j=1}^{m}\sum_{t=1}^{T}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq\omega_{j},\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right],$
@@
$G_{j}^{t}:=\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{i\in\{\pm 1\}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq i,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right],$ (K.33)
where we group $\bm{\omega}_{[-j]}^{1}$ and $\bm{\omega}_{[-j]}^{-1}$ together in the inner sum. Taking Equation (K.33) into Equation (K.32), we have
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)\gtrsim 2^{-m}M^{-\beta}\sum_{j=1}^{m}\sum_{t=1}^{T}G_{j}^{t}.$ (K.34)
We pause to provide some intuition for introducing $G_{j}^{t}$. The idea is that we would like to apply Bretagnolle-Huber inequality as stated in Theorem K.1 to obtain a lower bound of the cumulative regret. To get a tighter lower bound, we would group the most similar pairs of $\bm{\omega},\bm{\omega}^{\prime}\in\Omega_{m}$ together to minimize the KL divergence between the two probability measures indexed by $\bm{\omega}$ and $\bm{\omega}^{\prime}$.
By Equation (K.25) and the definition of $\delta_{0}$ in Equation (K.28),
$\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq i,\bm{O}_{t}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{t-1}\right]$ $\asymp\frac{1}{M^{d_{\texttt{non}}}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq i\right\}\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right].$ (K.35)
Denote by $\mathbb{P}^{(t-1)}_{j}$ the conditional probability $\mathbb{P}\left(\cdot\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right)$. We apply Bretagnolle-Huber inequality as stated in Theorem K.1 and obtain
$\sum_{i\in\{\pm 1\}}\mathbb{E}^{(t-1)}_{\bm{\theta},f_{\bm{\omega}^{i}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{t}\neq i\right\}\mid\mathcal{H}_{t-1},\bm{O}_{t}\in\widetilde{B}_{j}\right]$ (K.36)
@@
where the last inequality follows from Equations (K.16), (K.1) and (K.2). Taking Equations (K.35) and (K.36) into Equation (K.33) yields that
$G_{j}^{t}$ $\gtrsim M^{-d_{\texttt{non}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left[-\mathcal{K}_{\bm{\theta},t}^{(\texttt{non})}\left(f_{\bm{\omega}^{1}_{[-j]}},f_{\bm{\omega}^{-1}_{[-j]}}\right)\right]$ (K.37)
$\stackrel{{\scriptstyle(i)}}{{\geq}}\frac{1}{M^{d_{\texttt{non}}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left(-\frac{C}{M^{2\beta}}\sum_{\tau=1}^{t}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(t+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
$\geq\frac{1}{M^{d_{\texttt{non}}}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\exp\left(-\frac{C}{M^{2\beta}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(T+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}\frac{2^{m-1}}{M^{d_{\texttt{non}}}}\exp\left(-\frac{C}{M^{2\beta}2^{m-1}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}=-1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]-\frac{C(T+N)}{M^{2\beta+d_{\texttt{non}}}}\right)$
where inequality $(i)$ follows from Equation (K.31) and inequality $(ii)$ follows from Jensen’s inequality. Let
$E_{j,\pi}:=\frac{1}{2^{m-1}}\sum_{\bm{\omega}_{[-j]}\in\Omega_{m-1}}\sum_{\tau=1}^{T}\mathbb{E}^{(\tau-1)}_{\bm{\theta},f^{1}_{\bm{\omega}_{[-j]}},\pi}\mathbb{E}\left[\mathds{1}\left\{A_{\tau}=1,\bm{O}_{\tau}\in\widetilde{B}_{j}\right\}\mid\mathcal{H}_{\tau-1}\right]$
and taking $E_{j,\pi}$ into Equation (K.37), we have
$G_{j}^{t}\gtrsim\frac{2^{m-1}}{M^{d_{\texttt{non}}}}\exp\left(-CM^{-2\beta}E_{j,\pi}-C(T+N)M^{-2\beta-d_{\texttt{non}}}\right)$ (K.38)
From the definition of $G_{j}^{t}$ in Equation (K.33), we also have
$\sum_{t=1}^{T}G_{j}^{t}\geq 2^{m-1}E_{j,\pi}.$ (K.39)
Taking Equations (K.38) and (K.39) into Equation (K.34) yields
@@
$\sup_{f\in\mathcal{F}_{\beta,L}}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)$ $\gtrsim\inf_{z\geq 0}mM^{-\beta}\left\{z+\frac{T}{M^{d_{\texttt{non}}}}\exp\left[-CM^{-2\beta}z-\frac{CN}{M^{2\beta+d_{\texttt{non}}}}\right]\right\}$ (K.40)
@@
where in the last inequality, we use the definition of $m$ as in Equation (K.9) and let
@@
The minimizer of the right-hand side of Equation (K.40) over $z\in{{\mathbb{R}}}$ is given by
$z^{\star}=\frac{M^{2\beta}}{C}\log\left(\frac{CT\alpha}{M^{2\beta+d_{\texttt{non}}}}\right)=\frac{M^{2\beta}}{C}\log\left(\frac{CT}{M^{2\beta+d_{\texttt{non}}}}\right)-\frac{N}{M^{d_{\texttt{non}}}}.$ (K.41)
@@
$M^{2\beta+d_{\texttt{non}}}\log\left(\frac{CT}{M^{2\beta+d_{\texttt{non}}}}\right)\geq CN.$ (K.42)
Noting that when $M^{2\beta+d_{\texttt{non}}}>CT$, the left hand side of the above display is negative. Thus, for Equation (K.42) to hold, we must have $M^{2\beta+d_{\texttt{non}}}=O(T)$, implying that
@@
$M^{2\beta+d_{\texttt{non}}}=\frac{CT}{e}.$ (K.43)
When $T\geq CN$ for some constant $C$ sufficiently large, $z^{\star}\geq 0$ holds. Taking $z=z^{\star}$ in Equation (K.40) yields
@@
where the last equality holds from the choice of $M$ in Equation (K.43).
@@
and is monotonically increasing when $z>z_{0}$, it follows that the minimizer of $h(z)$ over $z\geq 0$ when $z_{0}<0$ is attained at $z=0$. Comparing the form of the right-hand side of Equation (K.40) to $h(z)$ defined above yields that the minimizer is attained at $z=0$ and
$\sup_{f}\mathcal{R}^{(\texttt{non})}_{T}(\bm{\theta},f)\geq\frac{T}{M^{\beta}}\exp\left(-\frac{CN}{M^{2\beta+d_{\texttt{non}}}}\right).$ (K.44)
@@
$M=\left[\frac{C(2\beta+d_{\texttt{non}})N}{\beta}\right]^{\frac{1}{2\beta+d_{\texttt{non}}}}=\Theta\left(N^{\frac{1}{2\beta+d_{\texttt{non}}}}\right).$ (K.45)
Taking Equation (K.45) into the right-hand side of Equation (K.44) yields the desired result in Equation (K.13). ∎
Appendix L Derivation of Equivalent Formulations for UCB Exploration in Algorithm 1